The following information is provided about an open economy with a government. Use the
information to answer the questions that follow:
C = 450 + 0.4Y
I = 350
G = 150
X = 70
Z = 35 + 0.1Y
T = 0.15Y
Yf = 1550
Q.2.1 Calculate the level of autonomous spending in this economy. (2)
Q.2.2 Calculate the size of the multiplier
1
Expert's answer
2020-11-27T00:36:13-0500

Q.2.1 The level of autonomous spending in this economy is 450 + 350 + 150 + 70 - 35 = 985.

Q.2.2 The size of the multiplier is:

"m = \frac {1} {1 - (0.4×(1 - 0.15) - 0.1)} = 1.32."